# Optimizing Symbol Visibility through Displacement

### v2

@@ Abstract
In information visualization, the position of symbols often encodes associated data values. When visualizing data elements with both a numerical and a categorical dimension, positioning in the categorical axis admits some flexibility. This flexibility can be exploited to reduce symbol overlap, and thereby increase legibility. In this paper we initialize the algorithmic study of optimizing symbol legibility via a limited displacement of the symbols.
Specifically, we consider unit square symbols that need to be placed at specified $y$-coordinates. We optimize the drawing order of the symbols as well as their $x$-displacement, constrained within a rectangular container, to maximize the minimum visible perimeter over all squares. If the container has width and height at most $2$, there is a point that stabs all squares. In this case, we prove that a staircase layout is arbitrarily close to optimality and can be computed in $O(n\log n)$ time. If the width is at most $2$, there is a vertical line that stabs all squares, and in this case, we give a 2-approximation algorithm (assuming fixed container height) that runs in $O(n\log n)$ time. As a minimum visible perimeter of 2 is always trivially achievable, we measure this approximation with respect to the visible perimeter exceeding 2. We show that, despite its simplicity, the algorithm gives asymptotically optimal results for certain instances.
@@ 1 Introduction
In both examples above, the position of the symbol is fixed and cannot be changed without severely distorting the information it encodes. In other settings, such as map labeling, the position of a symbol is not completely fixed, instead the symbol (the label) needs to be placed in contact with a particular point on the map. There are infinitely many potential placements for the symbol, but all must contain the same fixed point somewhere on its boundary. In this paper we consider a related symbol placement problem, which is motivated by the visualization of numerical data with associated categories: the age of employees within a certain division of a company or the page rank of tweets that exhibit a certain sentiment (positive, negative, neutral) on a topic such as vaccinations. Such data can be visualized in categorical strips of fixed width $w$, restricting the symbols to lie in the strip, and placing the symbols on a $y$-coordinate according to their numerical values (see Figure 1 for an example using twitter data). There are again infinitely many potential placements per symbol, but all placements are restricted to share the same $y$-coordinate.
@@
If the positions of symbols are fixed or restricted, then close symbols will overlap, reducing the visible part of – or even fully obscuring – other symbols. Correspondingly, there is an ample body of work on optimizing the visibility of symbols under placement restrictions. The algorithmic literature considers a couple of variants. First of all, we either display all symbols or only a subset. For symbol maps and our categorical strips we always have to display all symbols, since otherwise not all data is visualized. For map labeling one usually chooses a subset of the labels which can be placed without any overlap; the corresponding optimization problems attempt to maximize the number of these labels while also taking priorities (such as city sizes) into account. If overlap between symbols might be unavoidable, visibility is optimized by either maximizing the minimum perimeter of the symbol that has least visibility or maximising the total visible perimeter. If the positions of the symbols are completely fixed, then the only choice we can make is the drawing order of the symbols.
@@
Figure 1: A pro and con vaccination twitter discussion from $2018$, capturing $823$ accounts. Accounts are placed in a categorical strip according to most frequent sentiment expressed. Color indicates if account mostly mentions accounts with a similar opinion (blue) or with a different opinion (red). Top: random jittering, bottom: our approximation algorithm described in Theorem 4.1. Random jittering hides many details, such as pro-vaccination accounts that mention predominantly anti-vaccination accounts (red boundaries). Many accounts at the bottom sent the same number of tweets and hence share a $y$-coordinate, which inevitably covers most of their horizontal edges.
@@
In this paper we study the novel algorithmic question of how to optimize the visibility of a given set of symbols, all of which must be drawn, when we may choose their drawing order and their $x$-coordinate, given a set of fixed (and distinct) $y$-coordinates for each symbol. We measure the visibility of the result via the minimal visibility perimeter over all symbols [3]. Figure 1 shows that our algorithms do indeed greatly improve the visible perimeter and thereby give the viewer a more accurate impression of the data. Note that our theoretical results hold only for square symbols, but, as evidenced by Figure 1, the algorithms we propose readily extend to rectangular symbols. Proving similar bounds for more general symbol shapes is a challenging open problem.
@@
In this paper we initiate the algorithmic study of optimizing symbol visibility through displacement. Specifically, we focus on unit square symbols, that may be shifted horizontally while remaining in a strip of width 2 (their categorical strip). In Section 2 we introduce our notation and make some initial observations. Most notably: the visible perimeter behaves non-continuously when squares are placed on the same $x$-coordinate. Hence the optimal visible perimeter is a supremum that cannot always be reached. In Section 3 we study the special case that the strip has height at most 2. In this scenario all squares are stabbed by a point. We first establish several useful geometric properties of so-called reasonable layouts – arrangements of the input squares which meet certain lower bound conditions – and then use these properties to prove that a simple $O(n\log n)$ algorithm suffices to compute a layout of the squares whose visible perimeter is arbitrarily close to the supremum.
In Section 4 we then study the general case of strips of arbitrary height (but still width 2). Here all squares are stabbed by a line. We leverage our previous result to obtain an $O(n\log n)$-time approximation algorithm. This approximation is with respect to the so-called gap – the visible perimeter minus two – since a minimal visible perimeter of 2 is trivially obtained for any instance. Furthermore, if the $y$-coordinates are uniformly distributed, then we can show that a specific layout – the zigzag layout – is asymptotically optimal. We close with a discussion of various avenues for future work, both towards the practical applicability of our results in visualization systems and towards more theoretical results in other settings, including different visibility definitions and other symbol shapes.
@@
The questions we study in this paper combine various aspects and restrictions of well-known placement problems in the algorithmic (geo-)visualization literature in a novel way. Most closely related to our work are the two papers by Cabello et al. [3] and by Nivasch et al. [15] that consider perimeter problems for sets of differently sized symbols at fixed positions in the plane. Specifically, they consider the problems of maximizing the minimum perimeter of the symbol that has least visibility (as we do in this paper) or maximizing the total visible perimeter. Since the locations of the symbols are fixed, the algorithmic problem reduces to finding the optimal (not necessarily stacking) order of the symbols. This contrasts with our work where a limited form of displacement is allowed.
Labeling cartographic maps, where a subset of the labels are chosen such that they can be placed without any overlap, and the corresponding optimization problems are computationally complex in various settings [8], as they relate to maximum independent set. Constrained displacement of labels is often also allowed, and various of such placement models have been studied algorithmically [2, 4, 16, 17, 21]. The goal is always to place as many labels as possible without overlap. Displacing labels to the boundary of a map has attracted considerable algorithmic attention under different models as well, see [1] for a survey. However, such practice relies on leader lines to connect labels to the points being labeled; making it harder to identify data patterns and thus less suitable for visualizing data.
@@ 2 Preliminaries
Our input is a sequence of distinct $y$-coordinates $\mathbf{y}=(y_{1},y_{2},\ldots,y_{n})$. We want to find $x$-coordinates $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$, determining unit squares $s_{1},s_{2},\ldots,s_{n}$, where $s_{i}$ has centroid $(x_{i},y_{i})$. We also want to find a stacking order, so that the minimum visible perimeter among all squares is maximized.
More specifically, we are given a strip $T$ of width $w>1$ and height $h>1$, where we assume that $T=[0,w]\times[0,h]$, and that $\frac{1}{2}\leq y_{1}<y_{2}<\ldots<y_{n}\leq h-\frac{1}{2}$. Thus, in terms of the $y$-coordinates, we can speak about the highest, or the lowest, among any set of squares. We require $x_{i}\in[\frac{1}{2},w-\frac{1}{2}]$ for all $i$, so that all squares are in the strip. We say that $s_{i}$ is left (right) of $s_{j}$ if $x_{i}<x_{j}$ ($x_{i}>x_{j}$). This determines a leftmost and a rightmost square among any set of squares (ties handled as needed).
@@
The visible perimeter of a square in a layout is the total length of all its visible boundary, where a point on the boundary is visible if any other square $t$ containing it is behind $s$. The top square has visible perimeter $4$. The visible boundary is made up of (horizontal or vertical) visible edges. Note that each side of $s$ can host at most one visible edge, due to all squares having the same size.
The gap of a square in a layout is its visible perimeter minus $2$. If this is non-positive, we say that the square has no gap. The gap of a layout is the minimum of the gaps of all squares. This definition is motivated by the fact that we can always achieve a positive gap by suitable “standard” layouts which we introduce below. Ideally, we want to find an optimal layout, one that has the maximum gap among all layouts. However, this may not exist: one can easily construct instances where the only candidates for optimal layouts have duplicate $x$-coordinates, but no such layouts can actually be optimal, refer to Figure 2 for one such instance. Therefore, we take the supremum gap over all layouts as the benchmark which we want to approximate as closely as possible.
Figure 2: An instance with $y_{2}-y_{1}=2\varepsilon,y_{3}-y_{2}=\epsilon$ and a strip of width $w=1+\varepsilon$, where $\varepsilon\leq 0.2$. From only looking at the two highest squares $s_{2},s_{3}$, we see that the gap of every layout is bounded from above by $2\varepsilon$, and to achieve this gap, $s_{2}$ and $s_{3}$ together need to span the full strip width. Given this, $s_{1}$ cannot be the bottom square, since then it has visible perimeter at most $1+5\varepsilon\leq 2$ and hence no gap (left); $s_{2}$ cannot be the bottom square, either, as this would limit its visible perimeter to at most $1+2\varepsilon$ (middle). So $s_{3}$ has to be the bottom square (right). Since $x_{1}=x_{2}$ would give the bottom square among $s_{1},s_{2}$ a visible perimeter of at most $1+4\varepsilon$, we need to have $x_{1}\neq x_{2}$, but this means that $s_{3}$ has gap less than $2\varepsilon$. Hence, a gap of $2\varepsilon$ is not achievable, but any smaller gap is (by the right layout, as $x_{1}-x_{2}\rightarrow 0$).
@@
Proof.
We build a staircase facing right and up as in Figure 3 (left). Then the left and lower sides of each square are completely visible, as well as parts of its right and upper side. This yields a positive gap. ∎
@@
Figure 3: A staircase (left), three generalized staircases (right).
@@ 3.1 Reasonable layouts
Assume otherwise and consider the two squares $s_{\ell}$ and $s_{r}$ defining the left and right sides of the bounding box. If one of them is above $s$ and the other one below, the situation is as in Figure 4 (left). Since $s_{\ell}$ and $s_{r}$ overlap in both $x$- and $y$-coordinate, $s$ has horizontal and vertical visible edges of total length at most $1$ each. Thus, the visible perimeter of $s$ is at most 2. In the other case, $s_{\ell}$ and $s_{r}$ are w.l.o.g. both above $s$ as in Figure 4 (right). Then we consider the square $s_{d}$ defining the bottom side of the bounding box; w.l.o.g. $s_{d}$ is left of $s$. In this case, $s_{r}$ and $s_{d}$ prove that $s$ has no gap.
Figure 4: The visible perimeter of bottom square $s$ is at most two if it is contained in the bounding box of the other squares.
@@
If $\mathcal{A}$ is unreasonable, there is nothing to prove. Otherwise, let $t_{1},t_{2},\ldots,t_{n}$ be the sequence of squares in stacking order, i.e. $t_{1}\prec t_{2}\prec\cdots\prec t_{n}$, and let the bounding box of $t_{i},t_{i+1},\ldots,t_{n}$ be denoted by $\tau_{i}$ for $1\leq i\leq n$. Since $\mathcal{A}$ is reasonable, $t_{i}$ “sticks out” of $\tau_{i+1}$ (i.e. it is not contained in it) by Proposition 3.1. There are two cases: $t_{i}$ is a “corner square” (Figure 5 left), or a “side square” (Figure 5 middle and right). Let $\Delta X_{i}$ and $\Delta Y_{i}$ quantify by how much $t_{i}$ sticks out, horizontally and vertically. For a side square, one of those numbers is 00.
@@
This means that, for $1\leq i<n$, the intervals corresponding to $\Delta X_{i}$ (as well as those corresponding to $\Delta Y_{i}$) span disjoint $x$-intervals (or $y$-intervals) that are also disjoint from the two intervals (one in each coordinate) of length one that is spanned by the top square. Hence,
@@
This upper bound on the gap is easily seen to be tight. {observation} There are instances of $n$ squares for which a staircase has gap $\frac{w+h-2}{n-1}$.
Proof 3.5.

### v3

@@ Abstract
In information visualization, the position of symbols often encodes associated data values. When visualizing data elements with both a numerical and a categorical dimension, positioning in the categorical axis admits some flexibility. This flexibility can be exploited to reduce symbol overlap, and thereby increase legibility. In this paper, we initialize the algorithmic study of optimizing symbol legibility via a limited displacement of the symbols.
Specifically, we consider closed unit square symbols that need to be placed at specified $y$-coordinates. We optimize the drawing order of the symbols as well as their $x$-displacement, constrained within a rectangular container, to maximize the minimum visible perimeter over all squares. If the container has width and height at most $2$, there is a point that stabs all squares. In this case, we prove that a staircase layout is arbitrarily close to optimality and can be computed in $O(n\log n)$ time. If the width is at most $2$, there is a vertical line that stabs all squares, and in this case, we design a 2-approximation algorithm (assuming fixed container height) that runs in $O(n\log n)$ time. As it turns out that a minimum visible perimeter of 2 is always achievable with a generic construction, we measure this approximation with respect to the visible perimeter exceeding 2. We show that, despite its simplicity, the algorithm gives asymptotically optimal results for certain instances.
@@ 1 Introduction
In both examples above, the position of the symbol is fixed and cannot be changed without severely distorting the information it encodes. In other settings, such as map labeling, the position of a symbol is not completely fixed, instead the symbol (the label) needs to be placed in contact with a particular point on the map. There are infinitely many potential placements for the symbol, but all must contain the same fixed point somewhere on its boundary. In this paper, we consider a related symbol placement problem, which is motivated by the visualization of numerical data with associated categories: the age of employees within a certain division of a company or the page rank of posts that exhibit a certain sentiment (positive, negative, neutral) on a topic such as vaccinations. Such data can be visualized in categorical strips of fixed width $w$, restricting the symbols to lie in the strip, and placing the symbols on a $y$-coordinate according to their numerical values (see Figure 1 for an example using Twitter data). There are again infinitely many potential placements per symbol, but all placements are restricted to share the same $y$-coordinate.
@@
If the positions of symbols are fixed or restricted, then close symbols will overlap, reducing the visible part of – or even fully obscuring – other symbols. Correspondingly, there is an ample body of work on optimizing the visibility of symbols under placement restrictions. The algorithmic literature considers a couple of variants. First of all, we either display all symbols or only a subset. For symbol maps and our categorical strips we always have to display all symbols, since otherwise not all data is visualized. For map labeling one usually chooses a subset of the labels which can be placed without any overlap; the corresponding optimization problems attempt to maximize the number of these labels while also taking priorities (such as city sizes) into account. If overlap between symbols is unavoidable, visibility is optimized by either maximizing the minimum perimeter of the symbol that has least visibility or maximizing the total visible perimeter. If the positions of the symbols are completely fixed, then the only choice we can make is the drawing order of the symbols.
@@
Figure 1: A pro and con vaccination discussion on Twitter from $2018$, capturing $823$ accounts. Accounts are placed in a categorical strip according to the most frequent sentiment expressed. Color indicates if an account mostly mentions accounts with a similar opinion (blue) or with a different opinion (red). Top: random jittering, bottom: our approximation algorithm described in Theorem 4.1. Random jittering hides many details, such as the pro-vaccination accounts that mention predominantly anti-vaccination accounts (red boundaries). Many accounts at the bottom sent the same number of posts and hence share a $y$-coordinate, which inevitably covers most of their horizontal edges.
@@
In this paper, we study the novel algorithmic question of how to optimize the visibility of a given set of symbols, all of which must be drawn, when we may choose their drawing order and their $x$-coordinate, given a set of fixed (and distinct) $y$-coordinates for each symbol. We measure the visibility of the result via the minimum visibility perimeter over all symbols [3]. The symbols model data of the same category to be used inside a visualization with multiple categories. Hence the $x$-coordinates are restricted to lie in a small range, to ensure visual cohesion. Specifically, we study the setting where all symbols overlap horizontally. Symbols with identical $y$-coordinates (data values) require special care, both from a visual aggregation and an algorithmic point of view. In this first study, we focus on the general case of distinct $y$-coordinates. Figure 1 shows that our algorithms do indeed greatly improve the visible perimeter and thereby give the viewer a more accurate impression of the data. Note that our theoretical results hold only for square symbols, but, as evidenced by Figure 1, the algorithms we propose readily extend to rectangular symbols and not necessarily distinct $y$-coordinates. Proving similar bounds for more general symbol shapes is a challenging open problem.
@@
We initiate the algorithmic study of optimizing symbol visibility through displacement. Specifically, we focus on unit square symbols, that may be shifted horizontally while remaining in a strip of width 2 (their categorical strip). In Section 2, we introduce our notation and make some initial observations. Most notably: the visible perimeter behaves non-continuously when squares are placed on the same $x$-coordinate. Hence the optimal visible perimeter is a supremum that cannot always be reached. In Section 3 we study the special case that the strip has height at most 2. In this scenario all squares are stabbed by a point. We first establish several useful geometric properties of so-called reasonable layouts – arrangements of the input squares which meet certain lower bound conditions – and then use these properties to prove that a simple $O(n\log n)$ algorithm suffices to compute a layout of the squares whose visible perimeter is arbitrarily close to the supremum.
In Section 4 we then study the general case of strips of arbitrary height (but still width 2). Here all squares are stabbed by a line. We leverage our previous result to obtain an $O(n\log n)$-time approximation algorithm. This approximation is with respect to the so-called gap – the visible perimeter minus two – since a minimum visible perimeter of 2 is trivially obtained for any instance. Furthermore, if the $y$-coordinates are uniformly distributed, then we can show that a specific layout – the zigzag layout – is asymptotically optimal. We close with a discussion of various avenues for future work, both towards the practical applicability of our results in visualization systems and towards more theoretical results in other settings, including different visibility definitions and other symbol shapes.
@@
The questions we study in this paper combine various aspects and restrictions of well-known placement problems in the algorithmic (geo-)visualization literature in a novel way. Most closely related to our work are the two papers by Cabello et al. [3] and by Nivasch et al. [15] that consider perimeter problems for sets of differently sized symbols at fixed positions in the plane. Specifically, they consider the problems of maximizing the perimeter of the symbol that has least visibility (as we do in this paper), or maximizing the total visible perimeter. Since the locations of the symbols are fixed, the algorithmic problem reduces to finding the optimal (not necessarily stacking) order of the symbols. This contrasts with our work where a limited form of displacement is allowed.
Labeling cartographic maps, where a subset of the labels are chosen such that they can be placed without any overlap, and the corresponding optimization problems are computationally complex in various settings [8], as they relate to maximum independent set. Constrained displacement of labels is often also allowed, and various of such placement models have been studied algorithmically [2, 4, 16, 17, 21]. The goal is generally to place as many labels as possible without overlap, although some variants exist that allow labels to be scaled until all labels can be placed. By design, our symbols overlap horizontally and often also vertically, and we cannot omit symbols; furthermore, scaling symbols is not an option while retaining legibility of the visualization. Displacing labels to the boundary of a map has attracted considerable algorithmic attention under different models as well, see the paper [1] for a survey. However, such practice relies on leader lines to connect labels to the points being labeled; making it harder to identify data patterns and thus less suitable for visualizing data.
@@ 2 Preliminaries
Our input is a sequence of distinct $y$-coordinates $\mathbf{y}=(y_{1},y_{2},\ldots,y_{n})$. We would like to find $x$-coordinates $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$, determining unit squares $s_{1},s_{2},\ldots,s_{n}$, where $s_{i}$ has centroid $(x_{i},y_{i})$. We also want to find a stacking order, so that the minimum visible perimeter among all squares is maximized.
More specifically, we are given a strip $T$ of width $1<w\leq 2$ and height $h>1$, where we assume that $T=[0,w]\times[0,h]$, and that $\frac{1}{2}\leq y_{1}<y_{2}<\ldots<y_{n}\leq h-\frac{1}{2}$. Thus, in terms of the $y$-coordinates, we can speak about the highest, or the lowest, among any set of squares. Furthermore, we say that $s_{i}$ is above (below) $s_{j}$, if $y_{i}>y_{j}$ ($y_{i}<y_{j}$). To ensure visual cohesion between the squares, we require $x_{i}\in[\frac{1}{2},w-\frac{1}{2}]$ for all $i$, so that all squares are in the strip. We say that $s_{i}$ is left (right) of $s_{j}$, if $x_{i}<x_{j}$ ($x_{i}>x_{j}$). This determines a leftmost and a rightmost square among any set of squares (ties handled as needed).
@@
The visible perimeter of a square $s$ in a layout is the total length of all its visible boundary, where a point on the boundary is visible if any other square $t$ containing it is behind $s$. The top square has visible perimeter $4$. The visible boundary is made up of (horizontal or vertical) visible edges. Note that each side of $s$ can host at most one visible edge, due to all squares having the same size.
The gap of a square in a layout is its visible perimeter minus $2$. If this is non-positive, we say that the square has no gap. The gap of a layout is the minimum of the gaps of all squares. This definition is motivated by the fact that we can always achieve a positive gap by suitable “standard” layouts which we introduce below. Ideally, we want to find an optimal layout, one that has the maximum gap among all layouts. However, this may not exist: one can easily construct instances where the only candidates for optimal layouts have duplicate $x$-coordinates, but no such layouts can actually be optimal.
One example of such an instance has three squares, with $y_{2}-y_{1}=2\varepsilon,y_{3}-y_{2}=\epsilon$ and a strip of width $w=1+\varepsilon$, where $\varepsilon\leq 0.2$, see Figure 2. To see that there is no optimal layout in this case, from only looking at the two highest squares $s_{2},s_{3}$, we see that the gap of every layout is bounded from above by $2\varepsilon$, and to achieve this gap, $s_{2}$ and $s_{3}$ together need to span the full strip width. Given this, $s_{1}$ cannot be the bottom square, since then it has visible perimeter at most $1+5\varepsilon\leq 2$ and hence no gap (Figure 2, left); $s_{2}$ cannot be the bottom square, either, as this would limit its visible perimeter to at most $1+2\varepsilon$ (Figure 2, middle). So $s_{3}$ has to be the bottom square (Figure 2, right). Since $x_{1}=x_{2}$ would give the bottom square among $s_{1},s_{2}$ a visible perimeter of at most $1+4\varepsilon$, we need to have $x_{1}\neq x_{2}$, but this means that $s_{3}$ has gap less than $2\varepsilon$. Hence, a gap of $2\varepsilon$ is not achievable, but any smaller gap is (by the right layout, as $x_{1}-x_{2}\rightarrow 0$).
Therefore, we take the supremum gap over all layouts as the benchmark which we want to approximate as closely as possible.
Figure 2: An instance with $y_{2}-y_{1}=2\varepsilon,y_{3}-y_{2}=\epsilon$, and a strip of width $w=1+\varepsilon$, where $\varepsilon\leq 0.2$, has no optimal layout.
@@
Proof 2.2.
We build a staircase facing up and right, with no two squares sharing the $x$-coordinate, as in Figure 3 (far left). Then the left and lower sides of each square are completely visible, as well as parts of its right and upper side. This yields a positive gap.
@@
Figure 3: A staircase (far left), and three generalized staircases.
@@ 3.1 Reasonable layouts
Assume otherwise and consider some two squares $s_{\ell}$ and $s_{r}$ defining, respectively, the left and right sides of the bounding box. If one of them is above $s$ and the other one below, the situation is as in Figure 4 (left). Since $s_{\ell}$ and $s_{r}$ overlap in both $x$- and $y$-coordinate, $s$ has horizontal and vertical visible edges of total length at most $1$ each. Thus, the visible perimeter of $s$ is at most 2. In the other case, $s_{\ell}$ and $s_{r}$ are w.l.o.g. both above $s$ as in Figure 4 (right). Then we consider the square $s_{d}$ defining the bottom side of the bounding box; w.l.o.g. $s_{d}$ is left of $s$. In this case, $s_{r}$ and $s_{d}$ prove that $s$ has no gap.
Figure 4: The visible perimeter of the bottom square $s$ is at most two if it is contained in the bounding box of the other squares.
@@
If $\mathcal{A}$ is unreasonable, there is nothing to prove. Otherwise let $t_{1},t_{2},\ldots,t_{n}$ be the sequence of squares in stacking order, i.e. $t_{1}\prec t_{2}\prec\cdots\prec t_{n}$, and let the bounding box of $t_{i},t_{i+1},\ldots,t_{n}$ be denoted by $\tau_{i}$ for $1\leq i\leq n$. Since $\mathcal{A}$ is reasonable, $t_{i}$ “sticks out” of $\tau_{i+1}$ (i.e. it is not contained in it) by Proposition 3.1, which can be applied as a sublayout of a reasonable layout is clearly reasonable as well.
There are two cases: $t_{i}$ is a “corner square” (Figure 5 left), or a “side square” (Figure 5 middle and right). Let $\Delta X_{i}$ and $\Delta Y_{i}$ quantify by how much $t_{i}$ sticks out, horizontally and vertically. For a side square, one of those numbers is 00.
@@
This means that, for $1\leq i<n$, the (open) intervals corresponding to $\Delta X_{i}$ (as well as those corresponding to $\Delta Y_{i}$) span disjoint $x$-intervals (or $y$-intervals) that are also disjoint from the two intervals (one in each coordinate) of length one that is spanned by the top square. Hence,
@@
This upper bound on the gap is easily seen to be tight.
Lemma 3.5.
There are instances of $n$ squares for which a staircase has gap $\frac{w+h-2}{n-1}$.
Proof 3.6.